# Navigating the Social Welfare Frontier: Portfolios for Multi-objective Reinforcement Learning

### v1

@@ B.3 Healthcare Intervention
It is computationally intractable to optimally solve the RMAB problem [23]. Thus, we use the popular Whittle Index heuristic [35] to solve an RMAB problem. Specifically, the Whittle index quantifies the reward gain achieved for every arm in every state if we took the active action as compared to if we took the passive action. The policy then chooses the arms with the highest Whittle indices within the budget limit.
@@
B.4 Other Details
For Budget-ConstrainedPortfolio, we fix $p_{0}=-\frac{\ln N}{\ln(1/0.9)}\simeq 10.5\ln N$. The approximation factor $\mathcal{Q}(\Pi^{\prime})$ for any portfolio $\Pi^{\prime}$ is computed via a grid search over 1000 points in $(\infty,1]$ for the natural disaster environment and the healthcare intervention problem. For the taxi environment, we used 50 points due to the computational burden of solving each problem.

### v2

@@ E.3 Healthcare Intervention
It is computationally intractable to optimally solve the RMAB problem [Papadimitriou and Tsitsiklis, 1999]. Thus, we use the popular Whittle Index Heuristic [Whittle, 1988] to solve RMAB problem for a given single reward function. Specifically, Whittle Index quantifies the reward gain achieved for every arm in every state if we took the active action as compared to if we took the passive action. The policy then chooses the arms with highest whittle indices within the budget limit.
@@
E.4 Other Details
Initial $p$ for Budget-ConstrainedPortfolio
For Budget-ConstrainedPortfolio, we fix $p_{0}=-100$.
Computing actual approximation factor
The approximation factor $\mathcal{Q}(\Pi^{\prime})$ for any portfolio $\Pi^{\prime}$ is computed via a grid search over 1000 points in $(\infty,1]$ for the natural disaster environment and the healthcare intervention problem. For the taxi environment, we used 50 points due to the computational burden of solving each problem.
Choice of $\alpha$
For a given portfolio size $K$, we chose the smallest $\alpha\in\{0.05,0.10,0.15,\dots,0.90,0.95,0.99\}$ among which the resulting portfolio has size $K$.